How many ounces of an alloy containing 35% aluminum must be melted with an alloy containing 65% aluminum to obtain 30 ounces of an alloy containing 55% aluminum?
Answer by ankor@dixie-net.com(22740)

How many ounces of an alloy containing 35% aluminum must be melted with an alloy containing 65% aluminum to obtain 30 ounces of an alloy containing 55% aluminum?
:
Let x = amt of 35% alloy
then
(30-x) = amt of 65% alloy
:
A mixture equation
.35x + .65(30-x) = .55(30)
.35x + 19.5 - .65x = 16.5
.35x - .65x = 16.5 - 19.5
-.30x = -3
x = -3/-.3
x = +10 oz of the 35% alloy required